OpenStudy (anonymous):

I don't know what I'm doing wrong in solving this: (x-5)^3/x(x+2)<=0. I know you need to move x(x+2) to 0's side, but then I'm lost. Please and showing steps would be much obliged. :3

OpenStudy (alfie):

Is it like this? \[\frac{(x-5)^3} {x(x+2)} <= 0\] if so, just solve the numerator and get the solutions. Then solve the numerator. Then you have to merge the solutions into one graph.

OpenStudy (alfie):

You cannot bring the x(x+2) on the 0 side, since it is not an equation. You have to solve for \[(x-5)^3 <= 0\] AND \[x(x+2) <= 0\]

OpenStudy (anonymous):

oh I get it! okay so once you solve both sides do you add them together?

OpenStudy (anonymous):

so when you expand (x-5)^3 you would? I'm sorry, but I've never learned this.

OpenStudy (alfie):

No, you have to solve just the numerator, then get the solution. Put the solutions aside. Now, solve the denominator, get the solutions, put them aside. Now you have to put the solutions (both numerator and denominator) into a graph, so you get the solutions of the whole thing.

OpenStudy (anonymous):

what do you mean put into a graph?

OpenStudy (alfie):

Here's how you solve (a-b)^3 http://upload.wikimedia.org/wikipedia/it/math/1/7/8/178c646f7889d307d9bad19748ff0aec.png About the graph: When you make a graph such as this one: http://www3.wolframalpha.com/Calculate/MSP/MSP23919he76g64b031971000014208c2e9d9g6h70?MSPStoreType=image/gif&s=40&w=310&h=36 You just have to solve the first inequality, then the second, then merge the solutions.

OpenStudy (anonymous):

I can't find (x-5)^3's solution because I can't find the factor.

OpenStudy (alfie):

You can find it by:

OpenStudy (anonymous):

Thank you so much for all the help. :3 I realy haven't learned this in school yet, same with otherse in my class even though we are in Pre-cal. Thanks again!